# Comparison of renormalized interactions using one-dimensional few-body systems as a testbed

### v1

@@ B.1 Energy
$g\delta$ $0.9954\pm 2\cdot 10^{-4}$ $1.7566\pm 4\cdot 10^{-4}$ $0.4877\pm 3\cdot 10^{-4}$
$g(n)\delta$ $1.00151\pm 2\cdot 10^{-5}$ $1.836\pm 0.002$ $1.1802\pm 5\cdot 10^{-4}$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $0.9984\pm 2\cdot 10^{-4}$ $2.14\pm 0.01$ $1.063\pm 0.003$
@@
$g\delta$ $2.6124\pm 5\cdot 10^{-4}$ $0.511\pm 0.008$ $0.86\pm 0.01$
$g(n)\delta$ $2.6004\pm 3\cdot 10^{-4}$ $0.226\pm 0.005$ $1.41\pm 0.01$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $2.6006\pm 4\cdot 10^{-4}$ $0.2636\pm 5\cdot 10^{-4}$ $1.319\pm 0.009$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $2.12\pm 0.02$ $0.433\pm 0.007$ $0.38\pm 0.05$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ $2.19303\pm 4\cdot 10^{-5}$ $-51\pm 2$ $3.14\pm 0.02$
@@
$g(n)\delta$ $6.4331\pm 1\cdot 10^{-4}$ $1.18\pm 0.05$ $1.71\pm 0.02$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $6.4338\pm 1\cdot 10^{-4}$ $0.93\pm 0.02$ $1.5\pm 0.1$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ $6.4313\pm 5\cdot 10^{-4}$ $0.02\pm 0.02$ $1.0\pm 0.4$
@@
Let us now consider the convergence of the energy for the $1+3$ system computed with the bare contact interaction, $g\delta$. For attractive interactions, we observe that $\sigma<0.5$ and that $c$ deviates considerably from the results of the renormalized interactions. We conclude that the number of data points used in the fitting procedure is not sufficient – we need to perform calculations with larger values of $n$ to get an accurate estimate of the energy in the limit $n\to\infty$. For the repulsive $1+3$ system, we see that $\sigma\approx 1$, indicating again the importance of higher order terms in Eq. (36).
For the energies of the $1+3$ system calculated using the effective interaction in relative coordinates and the running coupling constant approach, we find $\sigma<0.5$ for attractive interactions. Note that the non-monotonic behavior shown in Fig. 3 for $V_{\mathrm{eff}}^{\mathrm{rel}}$ cannot be reproduced with Eq. (37). Therefore, we only fit the data for a few large values of $n$, $n=21-31$, in this case. Nevertheless, the resulting values of $c$ are in good agreement with that for $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$, whose use leads to by far the largest value of $\sigma$ and to the most stable fit. For repulsive interactions, we observe that the data based upon running coupling constant and $V_{\mathrm{eff}}^{\mathrm{rel}}$ converge faster than the bare contact interaction. For $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$, the value of $\sigma$ has a high uncertainty: This is because the results for the energy are already nearly converged, which means that there is only little variation in the fitted data. This can be also seen from the fact that $A\approx 0$.
@@ B.2 Density
$g\delta$ $1.4276\pm 3\cdot 10^{-4}$ $-0.4137\pm 7\cdot 10^{-4}$ $0.887\pm 0.001$
$g(n)\delta$ $1.4338\pm 4\cdot 10^{-4}$ $-0.83\pm 0.05$ $1.31\pm 0.03$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $1.43441\pm 6\cdot 10^{-5}$ $-0.72\pm 0.05$ $1.215\pm 0.003$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ $1.43607\pm 2\cdot 10^{-5}$ $-3.1\pm 0.1$ $2.47\pm 0.02$
@@
$g\delta$ $1.6001\pm 3\cdot 10^{-4}$ $-0.297\pm 0.005$ $0.85\pm 0.01$
$g(n)\delta$ $1.6069\pm 5\cdot 10^{-4}$ $-0.31\pm 0.03$ $1.10\pm 0.05$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $1.6076\pm 2\cdot 10^{-4}$ $-0.295\pm 0.006$ $1.02\pm$ 0.01
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ $1.6071\pm 4\cdot 10^{-4}$ $-0.04\pm 0.08$ $2\pm 1$
@@
As we have discussed in the main text, the behavior of the energy as a function of $n$ is not a herald of convergence properties for other observables. Therefore, we continue our analysis of convergence patterns by considering the density at $x=0$ in this subsection, see Tab. B 2. The conclusions we can draw are similar to the ones above. The results computed with renormalized interactions have larger values of $\sigma$, which demonstrates that not only are the results closer to the limit $n\to\infty$, they also converge faster. The results based upon $V_{\mathrm{eff}}^{\mathrm{rel}}$ and the running coupling approach are similar. The potential $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ has the strongest performance. The corresponding value of $\sigma$ is the largest of all. Further, the data are already nearly constant as can be seen by the fact that $A\approx 0$ within the error bars.
@@ B.3 One-body Hamiltonian
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $1.9527\pm 7\cdot 10^{-4}$ $-1.07\pm 0.01$ $0.879\pm 0.007$
@@
$g\delta$ $0.5919\pm 2\cdot 10^{-4}$ $-0.261\pm 0.006$ $1.00\pm 0.01$
$g(n)\delta$ $0.5883\pm 4\cdot 10^{-4}$ $-0.323\pm 0.004$ $0.726\pm 0.007$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.59698\pm 2\cdot 10^{-5}$ $-0.150\pm 0.006$ $1.56\pm 0.02$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.5887\pm 4\cdot 10^{-4}$ $-0.325\pm 0.003$ $0.716\pm 0.007$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $0.5834\pm 7\cdot 10^{-4}$ $-0.43\pm 0.01$ $0.86\pm 0.01$
@@
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $4.0187\pm 4\cdot 10^{-4}$ $56\pm 6$ $2.71\pm 0.04$
@@
$g\delta$ $2.5332\pm 8\cdot 10^{-4}$ $-4.0\pm 0.1$ $1.35\pm 0.01$
$g(n)\delta$ $2.5262\pm 8\cdot 10^{-4}$ $-1.99\pm 0.02$ $0.981\pm 0.007$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $2.5741\pm 6\cdot 10^{-4}$ $-1.7\pm 0.2$ $1.62\pm 0.05$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $2.575092\pm 9\cdot 10^{-6}$ $-60\pm 20$ $3.9\pm 0.1$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $2.5281\pm 8\cdot 10^{-4}$ $-1.92\pm 0.02$ $0.950\pm 0.006$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $2.530\pm 0.002$ $-1.99\pm 0.05$ $0.95\pm 0.01$
@@
$g\delta$ $0.31782\pm 9\cdot 10^{-5}$ $0.219\pm 0.002$ $0.863\pm 0.004$
$g(n)\delta$ $0.3153\pm 2\cdot 10^{-4}$ $0.480\pm 0.003$ $1.295\pm 0.003$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.31474\pm 2\cdot 10^{-5}$ $0.480\pm 0.001$ $1.191\pm 0.003$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.31515\pm 1\cdot 10^{-5}$ $0.459\pm 0.003$ $1.233\pm 0.001$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $0.31492\pm 2\cdot 10^{-5}$ $0.377\pm 0.007$ $1.672\pm 0.009$
@@
$g\delta$ $0.7399\pm 3\cdot 10^{-4}$ $0.302\pm 0.006$ $0.89\pm 0.01$
$g(n)\delta$ $0.73265\pm 1\cdot 10^{-5}$ $0.168\pm 0.005$ $1.66\pm 0.01$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.73277\pm 2\cdot 10^{-5}$ $0.153\pm 0.004$ $1.41\pm 0.01$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.73278\pm 2\cdot 10^{-5}$ $0.143\pm 0.004$ $1.42\pm 0.01$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $0.732442\pm 5\cdot 10^{-6}$ $0.088\pm 0.005$ $1.93\pm 0.03$
@@
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $2.070624\pm 7\cdot 10^{-6}$ $-3000\pm 800$ $5.13\pm 0.09$
@@
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $2.07192\pm 3\cdot 10^{-5}$ $8\cdot 10^{5}\pm 2\cdot 10^{5}$ $6.9\pm 0.3$
@@
$g\delta$ $3.0311\pm 8\cdot 10^{-4}$ $2.25\pm 0.06$ $1.20\pm 0.01$
$g(n)\delta$ $3.001\pm 0.001$ $0.9\pm 0.2$ $1.41\pm 0.09$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $3.001\pm 0.001$ $0.8\pm 0.1$ $1.33\pm 0.08$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $3.00227\pm 1\cdot 10^{-5}$ $1.6\pm 0.4$ $2.66\pm 0.09$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $3.001\pm 0.001$ $0.7\pm 0.1$ $1.30\pm 0.09$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $3.00241\pm 2\cdot 10^{-5}$ $2.7\pm 0.3$ $2.46\pm 0.04$
@@ B.4 Transition matrix elements
$g\delta$ $0.4367\pm 5\cdot 10^{-4}$ $-0.3490\pm 1\cdot 10^{-4}$ $0.377\pm 0.002$
$g(n)\delta$ $0.4283\pm 5\cdot 10^{-4}$ $-0.4348\pm 6\cdot 10^{-4}$ $0.449\pm 0.002$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.4202\pm 2\cdot 10^{-4}$ $-0.1642\pm 8\cdot 10^{-4}$ $0.983\pm 0.003$
@@
$g(n)\delta$ obs non-trans. $0.4181\pm 5\cdot 10^{-4}$ $-0.279\pm 0.001$ $0.544\pm 0.005$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.4271\pm 4\cdot 10^{-4}$ $-0.42149\pm 4\cdot 10^{-4}$ $0.448\pm 0.001$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.094920\pm 6\cdot 10^{-6}$ $0.0209\pm 4\cdot 10^{-4}$ $1.19\pm 0.01$
@@
$g(n)\delta$ obs non-trans. $0.0977\pm 1\cdot 10^{-4}$ $0.1034\pm 9\cdot 10^{-4}$ $0.695\pm 0.006$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.1014\pm 3\cdot 10^{-4}$ $0.263\pm 0.003$ $0.754\pm 0.006$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $0.1046\pm 4\cdot 10^{-4}$ $0.355\pm 0.009$ $0.89\pm 0.01$
@@
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $0.1518\pm 3\cdot 10^{-4}$ $47\pm 7$ $2.92\pm 0.07$
@@
$g\delta$ $0.1256\pm 2\cdot 10^{-4}$ $1.22\pm 0.03$ $1.34\pm 0.01$
$g(n)\delta$ $0.1265\pm 6\cdot 10^{-4}$ $0.77\pm 0.03$ $1.08\pm 0.02$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.1124\pm 3\cdot 10^{-4}$ $0.30\pm 0.07$ $1.4\pm 0.1$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $0.11239\pm 7\cdot 10^{-5}$ $0.2\pm 0.2$ $2.1\pm 0.4$
$g(n)\delta$ obs non-trans. $0.1198\pm 5\cdot 10^{-4}$ $0.50\pm 0.03$ $1.16\pm 0.03$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.2624\pm 6\cdot 10^{-4}$ $0.77\pm 0.03$ $1.06\pm 0.02$

### v2

@@ B.1 Energy
$g\delta$ $0.9954\pm 2\times 10^{-4}$ $1.7566\pm 4\times 10^{-4}$ $0.4877\pm 3\times 10^{-4}$
$g(n)\delta$ $1.00151\pm 2\times 10^{-5}$ $1.836\pm 0.002$ $1.1802\pm 5\times 10^{-4}$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $0.9984\pm 2\times 10^{-4}$ $2.14\pm 0.01$ $1.063\pm 0.003$
@@
$g\delta$ $2.6124\pm 5\times 10^{-4}$ $0.511\pm 0.008$ $0.86\pm 0.01$
$g(n)\delta$ $2.6004\pm 3\times 10^{-4}$ $0.226\pm 0.005$ $1.41\pm 0.01$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $2.6006\pm 4\times 10^{-4}$ $0.2636\pm 5\times 10^{-4}$ $1.319\pm 0.009$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $2.183\pm 0.007$ $0.74\pm 0.01$ $0.783\pm 0.009$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ $2.19341\pm 2\times 10^{-5}$ $-17.9\pm 0.1$ $2.76\pm 0.02$
@@
$g(n)\delta$ $6.4331\pm 1\times 10^{-4}$ $1.18\pm 0.05$ $1.71\pm 0.02$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $6.4338\pm 1\times 10^{-4}$ $0.93\pm 0.02$ $1.5\pm 0.1$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ $6.4313\pm 5\times 10^{-4}$ $0.02\pm 0.02$ $1.0\pm 0.4$
@@
Let us now consider the convergence of the energy for the $1+3$ system computed with the bare contact interaction, $g\delta$. For attractive interactions, we observe that $\sigma<0.5$. For the repulsive $1+3$ system, we see that $\sigma\approx 1$, indicating again the importance of higher order terms in Eq. (38).
For the energies of the $1+3$ system calculated using the effective interaction in relative coordinates and the running coupling constant approach, we find $\sigma\approx 0.78$ for attractive interactions. Note that the non-monotonic behavior shown in Fig. 3 for $g(n)\delta$ cannot be reproduced with Eq. (39). Therefore, we only fit the data for a few large values of $n$, $n=21-31$, in this case. To make the comparison with the other interactions faithful, we also use the restricted fitting regime for them. Nevertheless, the resulting values of $c$ are in good agreement with that for $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$, whose use leads to by far the largest value of $\sigma$ and to the most stable fit. For repulsive interactions, we observe that the data based upon running coupling constant and $V_{\mathrm{eff}}^{\mathrm{rel}}$ converge faster than the bare contact interaction. For $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$, the value of $\sigma$ has a high uncertainty: This is because the results for the energy are already nearly converged, which means that there is only little variation in the fitted data. This can be also seen from the fact that $A\approx 0$.
@@ B.2 Density
$g\delta$ $1.4276\pm 3\times 10^{-4}$ $-0.4137\pm 7\times 10^{-4}$ $0.887\pm 0.001$
$g(n)\delta$ $1.4338\pm 4\times 10^{-4}$ $-0.83\pm 0.05$ $1.31\pm 0.03$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $1.43441\pm 6\times 10^{-5}$ $-0.72\pm 0.05$ $1.215\pm 0.003$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ $1.43607\pm 2\times 10^{-5}$ $-3.1\pm 0.1$ $2.47\pm 0.02$
@@
$g\delta$ $1.6001\pm 3\times 10^{-4}$ $-0.297\pm 0.005$ $0.85\pm 0.01$
$g(n)\delta$ $1.6069\pm 5\times 10^{-4}$ $-0.31\pm 0.03$ $1.10\pm 0.05$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ $1.6076\pm 2\times 10^{-4}$ $-0.295\pm 0.006$ $1.02\pm$ 0.01
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ $1.6071\pm 4\times 10^{-4}$ $-0.04\pm 0.08$ $2\pm 1$
@@
As we have discussed in the main text, the behavior of the energy as a function of $n$ is not a herald of convergence properties for other observables. Therefore, we continue our analysis of convergence patterns by considering the density at $z=0$ in this subsection, see Tab. B 2. The conclusions we can draw are similar to the ones above. The results computed with renormalized interactions have larger values of $\sigma$, which demonstrates that not only are the results closer to the limit $n\to\infty$, they also converge faster. The results based upon $V_{\mathrm{eff}}^{\mathrm{rel}}$ and the running coupling approach are similar. The potential $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ has the strongest performance. The corresponding value of $\sigma$ is the largest of all. Further, the data are already nearly constant as can be seen by the fact that $A\approx 0$ within the error bars.
@@ B.3 One-body Hamiltonian
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $1.9527\pm 7\times 10^{-4}$ $-1.07\pm 0.01$ $0.879\pm 0.007$
@@
$g\delta$ $0.5919\pm 2\times 10^{-4}$ $-0.261\pm 0.006$ $1.00\pm 0.01$
$g(n)\delta$ $0.5883\pm 4\times 10^{-4}$ $-0.323\pm 0.004$ $0.726\pm 0.007$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.59698\pm 2\times 10^{-5}$ $-0.150\pm 0.006$ $1.56\pm 0.02$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.5887\pm 4\times 10^{-4}$ $-0.325\pm 0.003$ $0.716\pm 0.007$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $0.5834\pm 7\times 10^{-4}$ $-0.43\pm 0.01$ $0.86\pm 0.01$
@@
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $4.0187\pm 4\times 10^{-4}$ $56\pm 6$ $2.71\pm 0.04$
@@
$g\delta$ $2.539\pm 0.001$ $2.9\pm 0.1$ $1.22\pm 0.02$
$g(n)\delta$ $2.531\pm 0.001$ $-1.77\pm 0.05$ $0.93\pm 0.01$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $2.5723\pm 1\times 10^{-4}$ $-3.4\pm 0.2$ $1.92\pm 0.02$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $2.575092\pm 9\times 10^{-6}$ $-60\pm 20$ $3.9\pm 0.1$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $2.532\pm 0.002$ $-1.75\pm 0.05$ $0.90\pm 0.01$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $2.525\pm 0.001$ $-2.12\pm 0.06$ $0.99\pm 0.01$
@@
$g\delta$ $0.31782\pm 9\times 10^{-5}$ $0.219\pm 0.002$ $0.863\pm 0.004$
$g(n)\delta$ $0.3153\pm 2\times 10^{-4}$ $0.480\pm 0.003$ $1.295\pm 0.003$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.31474\pm 2\times 10^{-5}$ $0.480\pm 0.001$ $1.191\pm 0.003$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.31515\pm 1\times 10^{-5}$ $0.459\pm 0.003$ $1.233\pm 0.001$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $0.31492\pm 2\times 10^{-5}$ $0.377\pm 0.007$ $1.672\pm 0.009$
@@
$g\delta$ $0.7399\pm 3\times 10^{-4}$ $0.302\pm 0.006$ $0.89\pm 0.01$
$g(n)\delta$ $0.73265\pm 1\times 10^{-5}$ $0.168\pm 0.005$ $1.66\pm 0.01$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.73277\pm 2\times 10^{-5}$ $0.153\pm 0.004$ $1.41\pm 0.01$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.73278\pm 2\times 10^{-5}$ $0.143\pm 0.004$ $1.42\pm 0.01$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $0.732442\pm 5\times 10^{-6}$ $0.088\pm 0.005$ $1.93\pm 0.03$
@@
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $2.070624\pm 7\times 10^{-6}$ $-3000\pm 800$ $5.13\pm 0.09$
@@
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $2.07192\pm 3\times 10^{-5}$ $8\times 10^{5}\pm 2\times 10^{5}$ $6.9\pm 0.3$
@@
$g\delta$ $3.0307\pm 8\times 10^{-4}$ $2.16\pm 0.09$ $1.18\pm 0.02$
$g(n)\delta$ $3.00419\pm 4\times 10^{-5}$ $3.09\pm 0.07$ $1.918\pm 0.009$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $3.00227\pm 1\times 10^{-5}$ $2.21\pm 0.08$ $1.77\pm 0.02$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $3.00227\pm 1\times 10^{-5}$ $1.6\pm 0.4$ $2.66\pm 0.09$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $3.00460\pm 8\times 10^{-5}$ $2.21\pm 0.09$ $1.78\pm 0.02$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $3.00241\pm 2\times 10^{-5}$ $2.7\pm 0.3$ $2.46\pm 0.04$
@@ B.4 Transition matrix elements
$g\delta$ $0.4367\pm 5\times 10^{-4}$ $-0.3490\pm 1\times 10^{-4}$ $0.377\pm 0.002$
$g(n)\delta$ $0.4283\pm 5\times 10^{-4}$ $-0.4348\pm 6\times 10^{-4}$ $0.449\pm 0.002$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.4202\pm 2\times 10^{-4}$ $-0.1642\pm 8\times 10^{-4}$ $0.983\pm 0.003$
@@
$g(n)\delta$ obs non-trans. $0.4181\pm 5\times 10^{-4}$ $-0.279\pm 0.001$ $0.544\pm 0.005$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.4271\pm 4\times 10^{-4}$ $-0.42149\pm 4\times 10^{-4}$ $0.448\pm 0.001$
@@
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.094920\pm 6\times 10^{-6}$ $0.0209\pm 4\times 10^{-4}$ $1.19\pm 0.01$
@@
$g(n)\delta$ obs non-trans. $0.0977\pm 1\times 10^{-4}$ $0.1034\pm 9\times 10^{-4}$ $0.695\pm 0.006$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.1014\pm 3\times 10^{-4}$ $0.263\pm 0.003$ $0.754\pm 0.006$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs non-trans. $0.1046\pm 4\times 10^{-4}$ $0.355\pm 0.009$ $0.89\pm 0.01$
@@
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $0.1518\pm 3\times 10^{-4}$ $47\pm 7$ $2.92\pm 0.07$
@@
$g\delta$ $0.1256\pm 2\times 10^{-4}$ $1.22\pm 0.03$ $1.34\pm 0.01$
$g(n)\delta$ $0.1265\pm 6\times 10^{-4}$ $0.77\pm 0.03$ $1.08\pm 0.02$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs trans. $0.1124\pm 3\times 10^{-4}$ $0.30\pm 0.07$ $1.4\pm 0.1$
$V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ obs trans. $0.11239\pm 7\times 10^{-5}$ $0.2\pm 0.2$ $2.1\pm 0.4$
$g(n)\delta$ obs non-trans. $0.1198\pm 5\times 10^{-4}$ $0.50\pm 0.03$ $1.16\pm 0.03$
$V_{\mathrm{eff}}^{\mathrm{rel}}$ obs non-trans. $0.2624\pm 6\times 10^{-4}$ $0.77\pm 0.03$ $1.06\pm 0.02$